OpenStudy (anonymous):

Our sun in the Main Sequence(MS) burns 10% of its total mass in the core. In the nuclear fusion reaction 0.7% of mass is converted in to energy. The total energy output of the Sun in one second is L= 3.845x1026 J. The mass of the Sun M= 1.9891x1030 Kg. Using the Einstein’s famous equation for energy and mass, E=mc2 find the life time of the Sun in the main sequence in billions of years.

OpenStudy (anonymous):

m to be converted in energy: 10% * M * 0.7% total energy produced as MS: E=mc^2 Time it will take in seconds s=E/L s=(0.10 * 1.9891e30 Kg *0.007) * (3e8 m/s)^2 / 3.845e26 J

OpenStudy (anonymous):

s=3.26e17 s = 1.03e10 y Roughly speaking, 10 billion years

OpenStudy (anonymous):

i.e., since the sun is in this phase since roughly 5 billion years, still 5 to go.

OpenStudy (anonymous):

No need to rush for another planetary system, yet...

OpenStudy (anonymous):

Dimbeni I think this * mark says that the solar masses and all

OpenStudy (anonymous):

What means that s=E/S